# Fast Sampling of Protein Conformational Dynamics

### v2

@@ Abstract
Protein function does not solely depend on structure but often relies on dynamical transitions between distinct conformations. Despite this fact, our ability to characterize or predict protein dynamics is substantially less developed compared to state-of-the-art protein structure prediction. Molecular simulations provide unique opportunities to study protein dynamics, but the timescales associated with conformational changes generate substantial challenges. Enhanced sampling algorithms with collective variables can greatly reduce the computational cost of sampling slow processes. However, defining collective variables suitable to enhance sampling of protein conformational transitions is non-trivial. Low-frequency vibrations have long been considered as promising candidates for collective variable but their identification so far relied on assumptions inherently invalid at low frequencies. We recently introduced an analysis of molecular vibrations that does not rely on such approximations and remains accurate at low frequencies. Here, we modified this approach to efficiently isolate low-frequency vibrations in proteins and applied it to a set of five proteins of varying complexity. We demonstrate that our approach is not only highly reproducible but results in collective variables that consistently enhance sampling of protein conformational transitions and associated free energy surfaces on timescales compatible with high throughput applications. This enables the efficient generation of protein conformational ensembles, which will be key for future prediction algorithms aiming beyond static protein structures.
I Introduction
@@
Notably, this is not true for alternative approaches such as quasi-harmonic or principal component modes that aim to describe molecular vibrations based on a single co-variance matrix of atomic displacements. Correlations between pairs of low-frequency quasi-harmonic modes obtained from the same trajectories, as well as 2D and 3D sub-spaces spanned by two or three modes, rarely reach 0.5 (Figures S6 & S7 in the SI). Interestingly, extending the length of simulations to the microsecond timescale improves the reproducibility of co-variance matrix eigenvectors only moderately. This is shown in Figures S8 & S9 in the SI for principal component modes of HEWL computed from five 1 microsecond replica simulations. In contrast, the high reproducibility of low-frequency mode assignments with FRESEAN mode analysis demonstrates that, in addition to its ability to isolate anharmonic low-frequency vibrations [32], its underlying time-correlation formalism does not depend on rare events.
@@
The results are shown in Figure 3, where the extent of motion along the geometric variables is shown as box and whisker plots. A gray background indicates the range of motion consistently described by each of the five independent replicas. Blue and red symbols show the range of previously reported conformations for each system (for simplicity, we refer to them as "closed" and "open", respectively; see SI for details) and the insets display representative structures [42, 38, 41, 40, 39]. Both conformations and transitions between them are consistently explored in each simulation replica for HEWL, HIV-1 Protease and RBP. For MCL-1 and KRAS, we observed constant shifts between coordinates reported in the literature [39, 41] and free energy minima in our simulations (see next section). These shifts and free energy minima observed in our simulations are indicated as arrows and vertical dashed lines in Figure 3. Taking this into account, we find that for MCL-1 four out of five replicas sample both sates, while one replica (R5) does not reach the "open" state. For KRAS, we find that all five replicas sample the full range of T35-G60 distances characteristic of the "closed" to "open" transition, while two replicas (R1 and R5) do not fully sample G12-T35 distances characteristic for the "open" state. Thus considering all systems and five respective replicas, 22 out of 25 metadynamics trajectories (88%) sample transitions between previously reported states in less than 100 ns (using CVs extracted from unbiased 20 ns simulations). We note in passing that an extension of our metadynamics trajectories to $\approx$ 150 ns results in full sampling of "closed" to "open" transitions for every replica of each protein.
@@
A comparison to five replicas of unbiased 100 ns control simulations for each system in Figure S12 of the SI demonstrates the impact of our enhanced sampling scheme. For all systems, enhanced sampling along anharmonic low-frequency modes increased the consistently sampled range of motion at least two-fold (closer to sixfold for HIV-1 Protease and RBP). Despite the overall much more constricted dynamics in unbiased control simulations, individual trajectories for HEWL and KRAS sampled at least one full transition, indicating the unpredictability of rare events.
@@
FESs as a function of the respective CVs obtained directly from the five independent metadynamics simulations are shown in Figure S13 in the SI. The independently sampled CVs used in each simulation are similar as shown in Figure 1 but not identical, which hinders direct comparisons between replicas. However, after unbiasing the metadynamics trajectories, we projected weighted conformational ensembles on the geometric variables introduced in Figure 3 to construct FESs with a common set of variables [43]. In contrast to the biased distributions shown in Figure 3, the resulting FESs in Figure S14 in the SI allow us to distinguish high and low free energy conformations sampled during the metadynamics simulations. Correlations between geometric variables (sampled states fall on diagonal) indicate redundant information. We do not observe such correlations for the CVs used to enhance sampling in our simulations, i.e., modes 7-8 obtained from FRESEAN mode analysis at zero frequency, which are collective and orthogonal by definition.
@@
Comparisons of FESs obtained for distinct replica simulations of the same protein (Figure S14 in the SI) show that sampling remains insufficient to fully characterize the conformational ensembles. To overcome this limitation, we increased the number of replica metadynamics simulations for each system to 20 (starting structures sampled every 5 ns from unbiased 100 ns trajectories), each with 100 ns of metadynamics sampling and parallel processing. Here, we used identical vibrational modes as CVs (obtained for R1 above) in each replica to allow for direct comparisons and averaging of thermodynamic ensembles (see SI for details). In Figure 4, we plot the resulting FESs (obtained from averaged weighted probability distributions, see SI) both in the space defined by vibrational modes (panel A) and geometric variables (panel B). Statistical errors, i.e., the standard deviation describing the uncertainty of a single 100 ns simulation and the standard error of the mean for all 20 replica simulations, are shown as a function of the geometric variables in panel C. These statistical errors indicate uncertainties of less than $\pm$10 kJ/mol for a single 100 ns simulation and less than $\pm$3 kJ/mol after averaging. This is further emphasized in Figure S15 in the SI, in which we plot a 1-dimensional minimum free energy path for the "closed" to "open" transition in HEWL including statistical error bars for distinct numbers of trajectories. Such small uncertainties are remarkable given that error bars for protein FESs are often not even reported. [36, 38, 41, 40, 39] Converged FESs for five distinct protein systems obtained with the same general protocol highlights the role of anharmonic low-frequency vibrations as drivers of conformational change.
@@
We indicate previously reported conformations [42, 38, 41, 40, 39] (referred to as "closed" and "open" in analogy to Figure 3) with blue circles and red diamonds in panels B and C of Figure 4. Our simulations accurately reproduce previously reported conformations (apart from constant shifts for MCL-1 and KRAS), while revealing additional details on the FES that have so far not yet been described. The success of low-frequency vibrations to enhance sampling of previously observed conformational changes is readily explained upon projection of the "closed" and "open" state ensembles into our CV space. Notable, any pair of geometric variables (a single point in Figure 4B) describes an ensemble of conformations in alternative representations. In panel A of Figure 4, we indicate 40 conformations representative of the "closed" and "open" states for each system (selected via k-means clustering) in the CV spaces defined by low-frequency vibrations. With the exception of KRAS, the "closed" and "open" conformations are well separated in CV space (highlighted by ellipses), confirming that low-frequency vibrations used as CVs are directly related to the conformational dynamics reported previously in the literature.[36, 38, 40, 39] The lack of separation between "closed" and "open" states of KRAS in our CV space indicates that the low-frequency vibrations detected for KRAS describe independent collective motions that are not directly coupled to conformational dynamics reported previously for this system. Nevertheless, sampling is enhanced for KRAS due to indirect effects, e.g., the biased motion along vibrational modes allows the system to access alternative transition pathways with lower free energy barriers.
@@ Acknowledgements.
This work is supported by the National Science Foundation (CHE-2154834) and the National Institute of General Medical Sciences (R01GM148622). The authors acknowledge Research Computing at Arizona State University for providing high performance computing resources that have contributed to the research results reported within this work.
@@ I.1 Simulation Protocol
I.2 Well-Tempered Metadynamics Simulations
@@
I.3 Coarse-Grained Representation of Proteins
@@
I.4 All-Atom vs. Coarse-Grained Vibrational Density of States
@@
The all-atom VDoS describes all vibrations in the protein including bond vibrations etc. within the backbone and side chains that are most prominent in the vibrational fingerprint region. These vibrations have no impact on the COMs used for our coarse-grained representation. Thus the coarse-grained VDoS is primarily missing vibrational degrees of freedom at high frequencies while low-frequency vibrations corresponding to relative vibrations of secondary structure units, side chain motions etc. are largely preserved. Correspondingly, we find that 91% of the zero frequency motions are captured in the coarse-grained representation, while the information loss sets in gradually with increasing frequency.
@@
I.5 FREquency SElective ANharmonic (FRESEAN) Mode Analysis
@@
I.6 1D-VDoS for Eigenvector Projections
@@
To do so, we first project weighted velocities ($\tilde{\mathrm{v}}(t)$) from our simulation trajectories on a given eigenvector $\mathbf{Q}_{i}$ with its $3N$ components $q_{i}^{j}$.
@@
I.7 Reproducibility of FRESEAN Mode Analysis
@@
$C_{\mathrm{2D}}^{\left(k,l\right)}$ $=$ $\left(\sum_{i\in{7,8}}\sqrt{\sum_{j\in{7,8}}\left(Q_{i}^{(k)}\cdot Q_{i}^{(l)}\right)^{2}}\right)/2$ (S12)
$C_{\mathrm{3D}}^{\left(k,l\right)}$ $=$ $\left(\sum_{i\in{7,8,9}}\sqrt{\sum_{j\in{7,8,9}}\left(Q_{i}^{(k)}\cdot Q_{i}^{(l)}\right)^{2}}\right)/3$ (S13)
I.8 Quasi-Harmonic and Principal Component Modes
@@
I.9 Definition and Visualization of Geometric Variables
@@
Protein Geometric Variable Definition/Notes HEWL pincer angle angle between Cα-atom centers of 3 residue groups: 1. [28-31,111-114] 2. [90-93] 3. [44-45,51-52] radius of gyration radius of gyration computed with all protein atoms HIV-1 Pr flap distance distance between Cα-atoms of G51 in each monomer flap RMSD RMSD of residues I50-G52 in both monomers MCL-1 S255-T226 distance between Cα-atoms of residues S225 and T226 interdomain angle angle between Cα-atoms of residues S225, D241, T226 RBP twist angle dihedral angle between COMs of 4 groups of residues: 1. [1-100,236-259] 2. [99-100,236-237,258-259] 3. [108-109,230-231,269-270], 4. [108-231,269-271] hinge angle angle between COMs of 3 groups of residues: 1. [1-100,236-259] 2. [101-107,232-235,260-268] 3. [108-231,269-271] KRAS G12-T35 distance between Cα atoms of residues G12 and T35 T35-G60 distance between Cα atoms of residues T35 and G60
@@
I.10 "Closed" and "Open" Conformations
@@
I.11 Averaged Free Energy Surfaces
@@
I.12 Free Energy Surfaces in Distinct Variable Spaces
@@
I.13 Box & Whisker Plot Definition
@@
I.14 Unbiased Control Simulations
@@
I.15 Free Energy Surfaces for Individual Replicas in CV-Space
@@
I.16 Free Energy Surfaces for Individual Replicas in Geometric Space
@@
I.17 Minimum Free Energy Path for Conformational Transition in HEWL
Figure S15: To further visualize the statistical convergence of free energy surfaces obtained from 20 independent metadynamics simulations using low-frequency modes obtained from FRESEAN mode analysis as CVs, we re-plot in panel A the converged free energy surface for HEWL as a function of its geometric variables (same as in Figure 4 of the main text), while indicating the minimum free energy pathway connecting the "closed" and "open" conformations (dashed line). In panel B, we plot the average free energy profile along the minimum free energy path as a function of the pincer angle. Gray boxes are centered on the location of "closed" and "open" states as reported in the literature. These do not exactly correspond to the minima of our free energy surface, but fall within larger regions in which we find the FES to be approximately flat. In addition, we show error bars representing the standard deviation over our 20 simulations (red = statistical error of a single simulation) and the error of the mean obtained after averaging over 5, 10, 15 and all 20 simulations. Dashed horizontal lines indicate the uncertainty of the free energy at the minimum after averaging over all 20 simulations.

### v3

@@ Abstract
Protein function does not depend solely on structure but often relies on dynamical transitions between distinct conformations. Despite this fact, our ability to characterize or predict protein dynamics is substantially less developed compared to state-of-the-art protein structure prediction. Molecular simulations provide unique opportunities to study protein dynamics, but the timescales associated with conformational changes generate substantial challenges. Enhanced sampling algorithms with collective variables can greatly reduce the computational cost of sampling slow processes. However, defining collective variables suitable to enhance sampling of protein conformational transitions is nontrivial. Low-frequency vibrations have long been considered as promising candidates for collective variables, but their identification so far relied on assumptions inherently invalid at low frequencies. We recently introduced an analysis of molecular vibrations that does not rely on such approximations and remains accurate at low frequencies. Here, we modified this approach to efficiently isolate low-frequency vibrations in proteins and applied it to a set of five proteins of varying complexity. We demonstrate that our approach is not only highly reproducible, but results in collective variables that consistently enhance the sampling of protein conformational transitions and associated free energy surfaces on timescales compatible with high-throughput applications. This enables the efficient generation of protein conformational ensembles, which will be key for future prediction algorithms aiming beyond static protein structures.
Introduction
@@
Notably, this is not true for alternative approaches such as quasi-harmonic or principal component modes that aim to describe molecular vibrations based on a single co-variance matrix of atomic displacements. Correlations between pairs of low-frequency quasi-harmonic modes obtained from the same trajectories, as well as 2D and 3D sub-spaces spanned by two or three modes, rarely exceed 0.5 (Figures S6 & S7 in the SI). Interestingly, extending the length of simulations to the microsecond timescale improves the reproducibility of co-variance matrix eigenvectors only moderately. This is shown in Figures S8 & S9 in the SI for principal component modes of HEWL computed from five 1 microsecond replica simulations. In contrast, the high reproducibility of low-frequency mode assignments with FRESEAN mode analysis demonstrates that, in addition to its ability to isolate anharmonic low-frequency vibrations [32], its underlying time-correlation formalism does not depend on rare events.
@@
The results are shown in Figure 3, where the extent of motion along the geometric variables is shown as box and whisker plots. A gray background indicates the range of motion consistently described by each of the five independent replicas. Blue and red symbols show the range of previously reported conformations for each system (for simplicity, we refer to them as "closed" and "open", respectively; see SI for details) and the insets display representative structures [42, 38, 41, 40, 39]. Both conformations and transitions between them are consistently explored in each simulation replica for HEWL, HIV-1 Protease and RBP. For MCL-1 and KRAS, we observed constant shifts between coordinates reported in the literature [39, 41] and free energy minima in our simulations (see next section). These shifts and free energy minima observed in our simulations are indicated as arrows and vertical dashed lines in Figure 3. Taking this into account, we find that for MCL-1 four out of five replicas sample both sates, while one replica (R4) narrowly misses the shifted "open" state. For KRAS, we find that all five replicas sample the full range of T35-G60 distances associated with the "closed" to "open" transition (including the shifts). However, two replicas (R1 and R5) do not fully sample the G12-T35 distance of the shifted "open" state. Thus, considering all systems and five respective replicas, 22 out of 25 metadynamics trajectories (88%) sample known conformational transitions in less than 100 ns (using CVs extracted from unbiased 20 ns simulations). We note that extending the metadynamics trajectories to 155 ns results in full sampling of "closed" to "open" transitions for every replica of each protein.
@@
A comparison to five replicas of unbiased 100 ns control simulations for each system in Figure S12 of the SI demonstrates the impact of our enhanced sampling scheme. For all systems, enhanced sampling along anharmonic low-frequency modes increased the consistently sampled range of motion at least two-fold (closer to sixfold for HIV-1 Protease and RBP). Despite the overall much more constricted dynamics in the unbiased control simulations, individual trajectories for HEWL and KRAS sampled one full transition, indicating the unpredictability of rare events.
@@
FESs as a function of the respective CVs obtained directly from the five independent metadynamics simulations are shown in Figure S13 in the SI. The independently sampled CVs used in each simulation are similar as shown in Figure 1 but not identical, which hinders direct comparisons between replicas. However, after unbiasing the metadynamics trajectories, we projected weighted conformational ensembles on the geometric variables introduced in Figure 3 to construct FESs with a common set of variables [43]. In contrast to the biased distributions shown in Figure 3, the resulting FESs in Figure S14 in the SI allow us to distinguish high and low free energy conformations sampled during the metadynamics simulations. Correlations between geometric variables (sampled states fall on diagonal) indicate redundant information. Notably, we do not observe such correlations for the CVs used to enhance sampling in our simulations, i.e., modes 7-8 obtained from FRESEAN mode analysis at zero frequency, which are orthogonal collective motions by definition.
@@
Comparisons of FESs obtained for distinct replica simulations of the same protein (Figure S14 in the SI) show that sampling remains insufficient to fully characterize the conformational ensembles. To overcome this limitation, we increased the number of replica metadynamics simulations for each system to 20 (starting structures sampled every 5 ns from unbiased 100 ns trajectories), each with 100 ns of metadynamics sampling and parallel processing. Here, we used identical vibrational modes as CVs (obtained for R1 above) in each replica to allow for direct comparisons and averaging of thermodynamic ensembles (see SI for details). In Figure 4, we plot the resulting FESs (obtained from averaged weighted probability distributions, see SI) both in the space defined by vibrational modes (panel A) and geometric variables (panel B). Statistical errors, i.e., the standard deviation describing the uncertainty of a single 100 ns simulation and the standard error of the mean for all 20 replica simulations, are shown as a function of the geometric variables in panel C. These statistical errors indicate uncertainties of less than $\pm$10 kJ/mol for a single 100 ns simulation and less than $\pm$3 kJ/mol after averaging. This is further emphasized in Figure S15 in the SI, in which we plot a 1-dimensional minimum free energy path for the "closed" to "open" transition in HEWL including statistical error bars. Such small uncertainties are remarkable given that error bars for protein FESs are often not even reported. [36, 38, 41, 40, 39] The generation of converged FESs for five distinct protein systems, obtained with the same general protocol, highlights the role of anharmonic low-frequency vibrations as drivers of conformational change.
@@
We indicate previously reported conformations [42, 38, 41, 40, 39] (referred to as "closed" and "open" in analogy to Figure 3) with blue circles and red diamonds in panels B and C of Figure 4. Our simulations accurately reproduce previously reported conformations (apart from constant shifts for MCL-1 and KRAS), while revealing additional details on the FES that have so far not yet been described. The success of low-frequency vibrations to enhance sampling of previously observed conformational changes is readily explained upon projection of the "closed" and "open" state ensembles into our CV space. Notably, any pair of geometric variables (a single point in Figure 4B) describes an ensemble of conformations in alternative representations. In panel A of Figure 4, we indicate 40 conformations representative of the "closed" and "open" states for each system (selected by k-means clustering) in CV spaces defined by low-frequency vibrations. With the exception of KRAS, the "closed" and "open" conformations are well separated in CV space (highlighted by ellipses), confirming that the low-frequency vibrations used here as CVs are directly related to conformational dynamics reported previously for these proteins in the literature.[36, 38, 40, 39] The lack of separation between the "closed" and "open" states of KRAS in our CV space indicates that the low-frequency vibrations detected for KRAS describe independent collective motions that are not directly coupled to conformational dynamics reported previously for this system. Nevertheless, sampling is enhanced for KRAS due to indirect effects, e.g., the biased motion along vibrational modes allows the system to access alternative transition pathways with lower free energy barriers.
@@ Acknowledgements.
This work is supported by the National Science Foundation (CHE-2154834) and the National Institute of General Medical Sciences (R01GM148622). The authors acknowledge Research Computing at Arizona State University for providing high performance computing resources [50] that have contributed to the research results reported within this work.
@@
.2 Well-Tempered Metadynamics Simulations
@@
.3 Coarse-Grained Representation of Proteins
@@
.4 All-Atom vs. Coarse-Grained Vibrational Density of States
@@
The all-atom VDoS describes all vibrations in the protein including bond vibrations etc. within the backbone and side chains that are most prominent in the vibrational fingerprint region. These vibrations have no impact on the COMs used for our coarse-grained representation. Thus the coarse-grained VDoS is primarily missing vibrational degrees of freedom at high frequencies while low-frequency vibrations corresponding to relative vibrations of secondary structure units, side chain motions etc. are largely preserved. Correspondingly, we find that 88% of the zero frequency motions are captured in the coarse-grained representation, while the information loss sets in gradually with increasing frequency.
@@
.5 FREquency SElective ANharmonic (FRESEAN) Mode Analysis
@@
.6 1D-VDoS for Eigenvector Projections
@@
[…]
@@
.7 Reproducibility of FRESEAN Mode Analysis
@@
$C_{\mathrm{2D}}^{\left(k,l\right)}$ $=$ $\left(\sum_{i\in{7,8}}\sqrt{\sum_{j\in{7,8}}\left(Q_{i}^{(k)}\cdot Q_{j}^{(l)}\right)^{2}}\right)/2$ (S12)
$C_{\mathrm{3D}}^{\left(k,l\right)}$ $=$ $\left(\sum_{i\in{7,8,9}}\sqrt{\sum_{j\in{7,8,9}}\left(Q_{i}^{(k)}\cdot Q_{j}^{(l)}\right)^{2}}\right)/3$ (S13)
.8 Quasi-Harmonic and Principal Component Modes
@@
.9 Definition and Visualization of Geometric Variables
@@
Protein Geometric Variable Definition/Notes HEWL pincer angle angle between Cα-atom centers of 3 residue groups: 1. [28-31,111-114] 2. [90-93] 3. [44-45,51-52] radius of gyration radius of gyration computed with all protein atoms HIV-1 Pr flap distance distance between Cα-atoms of G51 in each monomer flap RMSD RMSD of residues I50-G52 in both monomers MCL-1 S255-T226 distance between Cα-atoms of residues S225 and T226 interdomain angle angle between Cα-atoms of residues S225, D241, T226 RBP twist angle dihedral angle between COMs of 4 groups of residues: 1. [1-100,236-259] 2. [99-100,236-237,258-259] 3. [108-109,230-231,269-270], 4. [108-231,269-271] hinge angle angle between COMs of 3 groups of residues: 1. [1-100,236-259] 2. [101-107,232-235,260-268] 3. [108-231,269-271] KRAS G12-T35 distance between Cα-atoms of residues G12 and T35 T35-G60 distance between Cα-atoms of residues T35 and G60
@@
.10 "Closed" and "Open" Conformations
@@
.11 Averaged Free Energy Surfaces
@@
.12 Free Energy Surfaces in Distinct Variable Spaces
@@
.13 Box & Whisker Plot Definition
@@
.14 Unbiased Control Simulations
@@
.15 Free Energy Surfaces for Individual Replicas in CV-Space
@@
.16 Free Energy Surfaces for Individual Replicas in Geometric Space
@@
.17 Minimum Free Energy Path for Conformational Transition in HEWL
Figure S15: To further visualize the statistical convergence of free energy surfaces obtained from 20 independent metadynamics simulations using low-frequency modes obtained from FRESEAN mode analysis as CVs, we re-plot in panel A the converged free energy surface for HEWL as a function of its geometric variables (same as in Figure 4 of the main text), while indicating the minimum free energy pathway connecting the "closed" and "open" conformations (dashed line). In panel B, we plot the average free energy profile along the minimum free energy path as a function of the pincer angle. Gray boxes are centered on the location of "closed" and "open" states as reported in the literature. These do not exactly correspond to the minima of our free energy surface, but fall within larger regions in which we find the FES to be approximately flat. In addition, we show error bars representing the standard deviation (red) and the error of the mean obtained after averaging over all 20 simulations (black).